Assume in the USA in the ’80s, the 25th percentile wage for the labor market overall remained stable, while there was a decrease in 50th percentile wage and an increase in 75th percentile wage. How would this affect the ratio of 25th percentile wage to 50th percentile wage, and the ratio of 50th percentile wage to 75th percentile wage?
Answer #1

Solution:

The question seems pretty much straightforward, we solve using simple mathematics.

Say 25th percentile wage =$x. This means that 25% of workers earned a wage below $x. This remained stable in the 80s.

Taking 50th percentile wage =$y. This means that 50% of workers earn a wage below $y, so this is also the median wage. Let the changed wage be $y', where $y' < $y.

Similarly, taking 75th percentile wage =$z, meaning 75% of workers earn a wage below $z. Let the changed wage be $z', where $z' > $z.

Thus, ratio of 25th percentile wage to 50th percentile wage:

Initially= x/y; new = x/y'. With same numerator and decreased denominator, this ratio increases.

In same way, ratio of 50th percentile wage to 75th percentile wage:

Initially, ratio = y/z, new ratio = y'/z'. With decreased numerator and increased numerator, this ratio decreases.